OpenStudy (anonymous):

write sin(cos-1x-tan-1y) in terms of x and y only where -1

OpenStudy (anonymous):

help. me

OpenStudy (anonymous):

Recall the identity, \[\sin(x\pm y)=\sin x\cos y\pm\sin y\cos x\]

OpenStudy (anonymous):

This means \[\sin(\cos^{-1}x-\tan^{-1}y)=\sin(\cos^{-1}x)\cos(\tan^{-1}y)-\sin(\tan^{-1}y)\cos(\cos^{-1}x)\]